# Analytics, have some humility: a statistical view of fourth-down decision making

### v4

@@ 4 Results
The traditional fourth-down decision procedure is based solely on effect size: make the decision that maximizes estimated win probability. This procedure ignores the uncertainty inherent in estimating win probability from noisy and highly correlated historical data. Thus, we modified the decision procedure to account for uncertainty: we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. Specifically, we measure our confidence in a decision by the proportion of bootstrapped win probability models that report the decision as optimal.
@@ 4.1 Example plays
We illustrate our decision procedure using example plays. Also, to compare our decision procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We detail the specification of this baseline coach model in Appendix H and include the model’s predictions in our decision figures.
@@
Example play 1. First, we compare Baldwin’s fourth-down decision making procedure to ours via example play 1 from Section 2.2. Baldwin views ${\mathsf{Go}}$ as a “strong” decision because he estimates that going for it provides a $3.6\%$ gain in win probability over attempting a field goal, shown in Figure 1. In Figure 4 we incorporate uncertainty quantification into decision making. Although our point estimate (the blue column) suggests that ${\mathsf{Go}}$ provides a $1.4\%$ gain in win probability over ${\mathsf{FG}}$, our $90\%$ confidence interval $[-3.9\%,5.1\%]$ suggests that ${\mathsf{Go}}$ could either be a terrible or a great decision. Also, $50\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column), reflecting massive uncertainty in the optimal fourth-down decision. So, we are not confident in the fourth-down recommendation provided by the point estimate.
Figure 4: Our decision chart for example play 1. The bootstrapped models are split amongst ${\mathsf{Go}}$ and ${\mathsf{FG}}$ (${\mathsf{boot\%}}$, in the orange column, is $50\%$-$50\%$), so we don’t have enough data to be confident in the fourth-down recommendation provided by the point estimate (blue column).
@@
Example play 2. Next, we compare Burke’s fourth-down decision procedure to ours via example play 2 from Section 2.2. Burke views ${\mathsf{Go}}$ as the right decision because the yellow dot in Figure 2 (representing the actual play’s yardline and yards to go) lies squarely in the red region and is far from the decision boundary. That chart, however, says nothing about the estimated strength of making the optimal decision or about uncertainty quantification. Hence, in Figure 5a we show our version of Burke’s chart, which uses color intensity to visualize the estimated gain in win probability by making a decision (darker colors indicate larger values).333 We use green for ${\mathsf{Go}}$, yellow for ${\mathsf{FG}}$, and red for ${\mathsf{Punt}}$ because we liked Burke’s Twitter thought that fourth-down decision making is like a traffic light. The pink dot (representing the actual play’s yardline and yards to go) lies in a moderately dark green region, indicating a moderate estimated gain in win probability by going for it. Being far from the decision boundary, however, does not imply it the best decision with certainty. Also, it is not obvious how far from the boundary is “far enough.” Hence in Figure 5b we provide an additional chart that illustrates uncertainty in the estimated optimal decision. Here, the color intensity indicates the proportion of bootstrapped models which make the estimated optimal decision. Much of the figure consists of dark colors, specifically dark green in the bottom and middle-left and dark red in the top and middle-right. For those combinations of yardline and yards to go, we are confident in the estimated optimal decision. For example play 2 (the pink dot), ${\mathsf{boot\%}}$ is high (about $90\%$), so we are confident that ${\mathsf{Go}}$ is the best decision. Other combinations of yardline and yards to go feature lighter colors, reflecting high uncertainty. For those game-states, we are not confident in the estimated optimal decision.
@@
Figure 5: Our decision boundary charts for example play 2. The left chart visualizes the estimated optimal decision (color) according to effect size (color intensity) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The pink dot (representing the actual play’s yardline and yards to go) lies squarely in the green region, so the point estimate says ${\mathsf{Go}}$. The right chart is similar except color intensity reflects bootstrap percentage. We are confident that ${\mathsf{Go}}$ is the best decision at this game-state (the pink dot) because it lies in a dark green region. Nonetheless, many other values of yardline and yards to go at this game-state are subject to large uncertainty (lighter color).
@@
We include more examples of our decision procedure in Appendix J. We also include an interactive Shiny App in which you can visualize our fourth-down decision procedure for any game-state on Github.444 https://github.com/snoopryan123/fourth_down/tree/main/3_shiny
4.2 Analytics, have some humility
Since win probability point estimates are subject to substantial uncertainty, we should not simply recommend the decision that maximizes estimated win probability. Due to the random idiosyncrasies in the history of football, which is contained in a relatively small dataset, win probability point estimates can make a decision seem strong even when it isn’t.
@@ 4.2 Analytics, have some humility
In Figure 6 we compare traditional fourth-down recommendations to our recommendations. The strength of a traditional fourth-down decision recommendation is proportional to effect size, the estimated gain in win probability by making that decision ($x$-axis). The strength of our fourth-down recommendation is proportional to the bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$). For all fourth-down plays from 2018-2022 in a given effect size bin, we plot the proportion of these decisions we view as confident, lean, and uncertain in Figure 6a. In Figure 6b we plot the frequency of these plays that lie in a given effect size bin.
@@
We are confident in the vast majority of decisions that have an effect size above $4\%$. Most decisions with an effect size under $1\%$ are uncertain, and decisions with an effect size between $1\%$ and $4\%$ are a healthy split between confident, lean, and uncertain. Many plays ($45\%$) have an effect size under $1\%$ and most plays ($87\%$) have an effect size under $4\%$. So, most fourth-down plays feature an effect size that is subject to considerable uncertainty. In particular, we are confident in just $34\%$ of all fourth-down decisions from 2018 to 2022 and a whopping $40\%$ of them are uncertain.
Refer to caption (a) Refer to caption (b)
Figure 6: Left: the proportion of observed fourth-down plays from 2018-2022 that lie in a given decision category (confident, lean, or uncertain according to ${\mathsf{boot\%}}$) as a function of estimated gain in win probability. Right: the frequency of observed fourth-down plays from 2018-2022 that lie in each bucket of estimated gain in win probability.
This analysis reflects substantial overconfidence in win probability point estimates; far fewer fourth-down decisions are as obvious as analysts claim. Although many decisions are estimated to provide solid gains in win probability, many of these perceived edges may actually be due to noise. Thus, we ask football analysts to have some humility: for many game-states there is simply not enough data to be confident in win probability point estimates.
@@
Further, we are confident in about $10\%$ of decisions with effect size under $1\%$ and in about $20\%$ of decisions with effect size between $1\%$ and $2\%$. Although decision confidence is correlated with estimated gain in win probability, they are fundamentally different: we can be confident in plays that provide small edges. Over the course of a season, these edges can accumulate into a large advantage that coaches should take advantage of.
Finally, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, in Appendix I we evaluate coaches only on plays for which we are confident.
@@
Although football analysts have been overconfident in their point estimates for many plays and have been underconfident in certain small edges for other plays, the football analytics community was largely correct that NFL coaches do not go for it enough on fourth down. Across all fourth-down plays from 2018-2022 that we are confident in, the coach made the right decision for $91\%$ of plays where they should have kicked but just $62\%$ of plays where they should have gone for it. Play calling in the NFL is still far too conservative: coaches consistently make wrong decisions, particularly when they should go for it.
@@ 5 Discussion
In this study we recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a fourth-down recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow that recommendation. Similarly, we believe it is arrogant for a football analyst to expect a coach to follow an uncertain fourth-down recommendation, regardless of the effect size. In today’s age in which sports analysts are fighting a public relations war to be taken seriously by owners and coaches, overconfidence in decision recommendations doesn’t help non-believers buy what we’re selling. Further, the coach is an on-field expert who spends a significant amount of time with his players and has access to information that doesn’t show up in the data (i.e., unobserved confounders). For instance, Eagles coach Nick Sirianni may notice that his usually dominant offensive line is missing a key player today, which does not currently appear in the data. When the estimated optimal fourth-down decision has too much uncertainty, we suggest leaving the decision to the coach’s discretion. Having said that, although football analysts have been overconfident in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
Although our analysis improves the state of the art, it is not without limitations and there are many avenues for future work. We discuss these in detail below and in Appendix K.
@@
Because statistical win probability models produce estimates that are too uncertain at many game-states, we suggest that future work explore probabilistic state-space models to estimate win probability. Probabilistic models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and win probability is calculated by simulating games. The effective sample size of transition probability models is the number of plays because they are fit from independent play-level observations. Some analysts in private industry have created proprietary probabilistic win probability models, which they believe are more accurate than statistical models because they have a larger effective sample size. The tradeoff is that state-space models have higher bias than statistical models. Whereas both classes of models have hidden bias, or bias due to not adjusting for unobserved confounders, just state-space models have model bias as they make simplifying assumptions.
Proprietors of probabilistic models believe that introducing bias in order to reduce variance improves the overall accuracy of the resulting win probability estimator. Nevertheless, these models are subject to their own set of issues, and we believe they aren’t as low-variance as some analysts claim. Properly modeling the distribution of the outcome of a play is nontrivial. In contrast to the simple binary win/loss outcome of statistical win probability models, the outcome variable of a transition probability model is the next game-state, which could include a change in yardline, score, time, or timeouts. This distribution is quite complex: there is a spike at gaining 0 yards for incompletions, a spike for a touchdown, spikes for penalties, and other smooth possibly multimodal distributions for the outcome of run or pass plays, each of which change as a function of team quality and other confounders. Typical transition probability models are riddled with selection bias, as a coach generally chooses play calls that work for his specific players and don’t generalize to an “average” team. Further, uncertainty in these transition probabilities may, after being properly propagated through a state-space model, result in just as much (if not more) uncertainty in estimated win probability than estimates from statistical models. Additionally, one must take great care to carefully encode all the subtle rules of football into her model, and one needs sufficient computing power to simulate enough games to estimate win probability with enough granularity. We look forward to a public facing exploration of probabilistic win probability models in the future.
@@ Appendix A Data details
score differential
point differential between the offensive and defensive
team at the start of this play
@@
yardline
num. yards from the opponent’s endzone in
$\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
ydstogo
the number of yards the offense has left to gain
before achieving a first down or a touchdown
@@
roof
categorical variable grouping the roof of the stadium into
@@
Appendix B Estimating the win probability of a fourth-down decision
In this Section, we estimate the win probability of a fourth-down decision in $\{{\mathsf{Punt}},{\mathsf{FG}},{\mathsf{Go}}\}$ in terms of first-down win probability (denoted $\mathsf{WP_{1}}$) and decision transition probabilities (punt outcome distribution, field goal success probability, and conversion probability). We estimate first-down win probability in Section 2.4 and decision transition probabilities in Appendix C.
Win probability of ${\mathsf{Punt}}$. Suppose the offensive team has a fourth down at yardline $y$ and denote the remainder of the game-state by ${\mathbf{x}}$. If the offensive team punts, the opposing team has a first down at the subsequent yardline, which we model as a random variable. Hence the win probability of punting is one minus the opponent’s first-down win probability at the next yardline $y^{\prime}$ and next game-state ${\mathbf{x}}^{\prime}$,
We model the expected next yardline after punting as a function of yardline and punter quality in Appendix C.1.
Win probability of ${\mathsf{FG}}$. We decompose field goal win probability in terms of field goal success probability and first-down win probability on the subsequent play,
[…]
We model the probability of a successful field goal as a function of yardline and kicker quality in Appendix C.2. If the kicking team misses the field goal and the spot of the kick was within the 20 yardline, the opposing team has a first down at yardline $80$. If the kicking team misses the field goal and the spot of the kick was beyond the 20 yardline, the opposing team has a first down at the spot of the kick. The spot of the kick is typically 7 yards behind the line of scrimmage, so in terms of the current yardline $y$, the next yardline relative to the opposing team is $\min\{80,100-(y+7)\}$. Denoting the remaining game-state variables relative to the opposing team at the next play by ${\mathbf{x}}^{\prime}$, we have
[…]
If the kicking team makes the field goal, it scores three points and the opposing team has a first down after a kickoff. Denoting the score differential relative to the kicking team by $s$, we have
${\mathsf{WP}}(\text{make FG})=1-{\mathbb{E}}_{\text{kickoff}}[\mathsf{WP_{1}}(\text{yardline }y^{\prime},s^{\prime}=-s-3,{\mathbf{x}}^{\prime})].$ (16)
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }{\mathbb{E}}_{\text{kickoff}}[y^{\prime}],\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (17)
${\mathsf{WP}}(\text{make FG})\approx 1-\mathsf{WP_{1}}(\text{yardline }75,\ s^{\prime}=-s-3,{\mathbf{x}}^{\prime}).$ (18)
Win probability of ${\mathsf{Go}}$. Suppose the offensive team has a fourth down and $z$ yards-to-go at yardline $y$. If the offensive team goes for it and gains $\Delta\geq z$ yards, then in the next play it either has a first down at yardline $y-\Delta$ or a touchdown. Conversely, if the offensive team goes for it and gains $\Delta<z$ yards, then in the next play the opposing team has a first down at yardline $100-(y-\Delta)$. Hence the expected value of going for it on fourth down is
$\sum_{\Delta\geq z}\mathsf{WP_{1}}(\text{yardline }y-\Delta)\cdot{\mathbb{P}}(\text{gain $\Delta$ yards})+\sum_{\Delta<z}(1-\mathsf{WP_{1}}(\text{yardline }100-(y-\Delta)))\cdot{\mathbb{P}}(\text{gain $\Delta$ yards)}.$) . (19)
Here, the first-down win probability and conversion success probability are implicitly functions of game-state. As the probability distribution of the yards gained on a conversion attempt is complex, we employ a standard simplification from Burke, 2009b . If the offensive team converts on fourth down and $z$ yards-to-go at yardline $y$, we assume they gain $z$ yards on that play, yielding a first down at yardline $y-z$ or a touchdown. If the offensive team fails to convert, we assume they turn the ball over at the initial yardline of the play, leaving the opposing team with a first down at yardline $100-y$. Thus, the win probability of going for it is
$\begin{split}\mathsf{WP_{1}}(\text{yardline }y-z,{\mathbf{x}}^{\prime})\cdot{\mathbb{P}}(\text{convert}|{\mathbf{x}})+(1-\mathsf{WP_{1}}(\text{yardline }100-y,{\mathbf{x}}^{\prime}))\cdot(1-{\mathbb{P}}(\text{convert}|{\mathbf{x}})),\end{split}$ (20)
where ${\mathbf{x}}$ is the game-state on fourth down and ${\mathbf{x}}^{\prime}$ is the game-state on the subsequent play (as before, ${\mathbf{x}}^{\prime}$ flips the game-state variables which are relative to the team with possession (e.g., score differential, team quality metrics, timeouts remaining) and doesn’t alter the other variables (e.g., time remaining)). We model conversion probability as a function of yardline, yards-to-go, and offensive and defensive quality in Appendix C.3.
Appendix C Estimating fourth-down decision transition probabilities
In this Section we estimate fourth-down decision transition probability models. In Section C.1 we model the expected next yardline after punting, in Section C.2 we model field goal success probability, and in Section C.3 we model conversion success probability.
C.1 Modeling punt outcome
In this Section we model the outcome of a punt. Specifically, we model the expected next yardline after a punt (from the perspective of the receiving team) as a function of yardline and punter quality (from the perspective of the punting team).

### v5

@@ 4 Results
In the previous section we modified the decision procedure to account for uncertainty. In Section 4.1 we illustrate this amended procedure using example plays. Then, in Section 4.2 we consider the extent to which the traditional decision procedure was overconfident in its fourth-down recommendations.
@@ 4.1 Example plays
Now, we discuss example plays. To compare decision recommendations to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We detail the specification of this baseline coach model in Appendix D and include the model’s predictions in our decision figures.
@@
Example play 1. First, recall example play 1 from Section 2.2 in which the Patriots had the ball against the Colts in Week 10 of 2023. In Figure 5 we knit uncertainty quantification into decision making. Although the point estimate (the blue column in Figure 5a) suggests that ${\mathsf{Go}}$ provides a $1.5\%$ gain in win probability over ${\mathsf{FG}}$, $40\%$ of the bootstrapped models view ${\mathsf{FG}}$ as better than ${\mathsf{Go}}$ (the orange column in Figure 5a), reflecting substantial uncertainty in the optimal fourth-down decision. Also, the $90\%$ confidence interval is $[-6\%,6\%]$, and the histogram of bootstrapped win probability gain estimates in Figure 5b is centered near zero and has a large spread. These suggest that ${\mathsf{Go}}$ could either be a good or a bad decision.
Figure 5: Decision charts for example play 1. Figure (a) shows the models’ outputs, including the point estimate (blue column), ${\mathsf{WP}}$ gain confidence interval, ${\mathsf{boot\%}}$ (orange column), and other metrics. Figure (b) is the histogram of bootstrapped win probability gain estimates. The gray dashed line is zero and the blue solid line is the point estimate.
@@
Example play 2. Next, recall example play 2 from Section 2.2 in which the Eagles had the ball against the 49ers in the 2023 NFC Championship game. We visualize fourth-down decision recommendations in Figure 6. The pink dot in Figure 6a lies in a moderately dark green region, indicating a moderate estimated gain in win probability by going for it. Being far from the decision boundary, however, does not imply it the best decision with certainty. Also, it is not obvious how far from the boundary is “far enough.” Hence, in Figure 6b we provide an additional chart that illustrates uncertainty in the estimated optimal decision. Here, the color intensity indicates the proportion of bootstrapped models that make the estimated optimal decision. Much of the figure consists of dark colors, specifically dark green in the bottom-left and middle and dark red in the top-right. For those combinations of yardline and yards to go, we are confident in the estimated optimal decision. For example play 2 (the pink dot), ${\mathsf{boot\%}}$ is high (dark green), so we are confident that ${\mathsf{Go}}$ is the best decision. Other combinations of yardline and yards to go feature lighter colors, reflecting higher uncertainty. For those game-states, we are less confident in the estimated optimal decision.
@@
Figure 6: Decision boundary charts for example play 2. Figure (a) visualizes the estimated optimal decision (color) according to effect size (color intensity, where darker colors indicate larger values) as a function of yardline ($x$-axis) and yards to go ($y$-axis), holding the other game-state variables constant. The pink dot represents the actual play’s yardline and yards to go. Figure (b) is similar except color intensity reflects bootstrap percentage.
@@
We include more examples of this fourth-down decision procedure in Appendix E. We also include an interactive Shiny App that visualizes the decision procedure for any game-state on Github.555 To run the Shiny app, download the Github repository at https://github.com/snoopryan123/fourth_down and run 3_shiny/app.R.
4.2 Quantifying overconfidence in win probability point estimates
@@
In Figure 7 we visualize the distributions of effect size and uncertainty across all fourth-down plays from 2018-2022. Effect size is the estimated gain in win probability and uncertainty is proportional to bootstrap percentage. To facilitate easier communication to a non-technical football coach, we bin decision confidence levels into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
Figure 7: The distribution of effect size, and uncertainty given effect size, across all fourth-down plays from 2018-2022. The height of a bar reflects the proportion of plays that lie in its associated effect size bin ($x$-axis). The color distribution within a bar reflects the distribution of decision confidence (confident in green, lean in yellow, and uncertain in red) among those plays.
@@
We are confident in the vast majority of decisions that have an effect size above $4\%$. Most decisions with an effect size under $1\%$ are uncertain, and decisions with an effect size between $1\%$ and $4\%$ are a healthy split between confident, lean, and uncertain. Many plays ($45\%$) have an effect size under $1\%$ and most plays ($87\%$) have an effect size under $4\%$. So, most fourth-down plays feature an effect size that is subject to considerable uncertainty. In particular, we are confident in just $30\%$ of all fourth-down decisions from 2018 to 2022 and $44\%$ of them are uncertain. This analysis reflects substantial overconfidence in win probability point estimates; far fewer fourth-down decisions are as obvious as traditional analyses suggest.
@@
Interestingly, we are confident in about $10\%$ of decisions with effect size under $1\%$ and in about $20\%$ of decisions with effect size between $1\%$ and $2\%$. Although decision confidence is correlated with estimated gain in win probability, they are fundamentally different: we can be confident in plays that provide small edges. Over the course of a season, these edges can accumulate into a large advantage that coaches should take advantage of.
Furthermore, a football analyst shouldn’t penalize a coach for making a decision that has high uncertainty regardless of the effect size because the estimated edge may be due to noise. Accordingly, we should evaluate coaches only on plays for which we are confident (and perhaps also plays in which we lean towards a decision). In Figure 8 we rank coaches by the proportion of fourth-down decisions they made in accordance with our model from 2018 to 2022 among fourth-down plays we are confident in. The top of this list has coaches like John Harbaugh from “analytics bent” organizations but also coaches who are considered more traditional on fourth down such as Ron Rivera, Sean McVay, and Mike Vrabel. While “analytics bent” coaches tend to follow fourth-down recommendations from win probability point estimates, which include recommendations on confident plays, some of the more traditional coaches seem to have a good feel for obvious fourth downs.
Figure 8: NFL coach ($y$-axis) versus the proportion of confident fourth-down decisions he made in agreement with our model from 2018 to 2022 ($x$-axis).
@@
Despite overconfidence in win probability point estimates, analysts have been largely correct that NFL coaches do not go for it enough on fourth down. Across all fourth-down plays from 2018-2022 that we are confident in, the coach made the right decision for $90\%$ of plays where they should have kicked but just $57\%$ of plays where they should have gone for it. Play calling in the NFL is still far too conservative: coaches consistently make wrong decisions, particularly when they should go for it.
@@ 5 Discussion
Win probability models are commonly used across sports analytics. Fourth-down decision making is a prominent example. There has been little reflection on the high-variance nature of win probability estimators and no substantial attempt to quantify uncertainty in these models. In this study, we argue it is imperative to knit uncertainty quantification into win probability-based decision making. Using fourth-downs as our case study, we in particular recommend a fourth-down decision when we are confident it has higher win probability than all other decisions. If we are not confident in a recommendation because it is subject to substantial uncertainty, we should not penalize a coach who doesn’t follow it. Similarly, we do not think it is right for a football analyst to recommend an uncertain fourth-down recommendation, regardless of the effect size. Despite analysts’ overconfidence in win probability point estimates, after accounting for uncertainty we still find that NFL coaches are too conservative: they don’t go for it enough on fourth down.
Our analysis is not without limitations and there are many avenues for future work. To begin, in this study, we used bootstrapping to quantify uncertainty in win probability estimates. Although our bootstrap procedure produces wide confidence intervals, it underestimates uncertainty since it quantifies sampling uncertainty but not model uncertainty. The former is uncertainty resulting from fitting a model on a finite dataset (“variance”) and the latter is uncertainty caused by our model being wrong or biased (“bias”). Our models are subject to several sources of model uncertainty. We made several simplyfing assumptions in modeling fourth-down win probability, such as swapping the expectation with first-down win probability. Furthermore, “true” win probability is a function of unobserved confounders. For example, how well Tom Brady slept the previous night could affect his team’s win probability. A more elaborate analysis would capture model uncertainty.
@@
In this study, we found that statistical win probability models produce uncertain estimates at many game-states. In future work we suggest exploring probabilistic state-space models to estimate win probability. As discussed in Section 2.4, probabilistic models simplify the game of football into a series of transitions between game-states. Transition probabilities are estimated from play-level data and win probability is calculated by simulating games. The effective sample size (${\mathrm{ESS}}$) of transition probability models is the number of plays because they are fit from independent play-level observations. Some analysts in industry have created proprietary probabilistic models, which they believe are more accurate than statistical models due to the larger ${\mathrm{ESS}}$. But, state-space models have higher bias than statistical models. Whereas both classes of models have similar levels of hidden bias (due to not adjusting for unobserved confounders), state-space models have more model bias as they make stronger simplifying assumptions. See Appendix F for a further discussion of the difficulty of formulating probabilistic state-space win probability models. We look forward to a public facing exploration of those models in the future.
Finally, if teams follow our fourth-down recommendations, their behavior will change and win probability will change accordingly. Statistical win probability models that learn from the game outcomes defining the recent history of football do not account for this distribution shift. State-space models, on the other hand, can account for these changes by altering the probability that a team goes for in on fourth down as a function of game-state. A more elaborate analysis would account for this distribution shift.
@@ Appendix A Data details
score differential point differential between the offensive and defensive team at the start of this play
@@
yardline num. yards from the opponent’s endzone in $\{0,1,...,99,100\}$ (00 is touchdown and $100$ is safety)
ydstogo the number of yards the offense has left to gain before achieving a first down or a touchdown
@@
Appendix B Estimating player / team quality